OpenStudy (anonymous):

4(a • b) = (4 • a)b is an example of which algebraic property?

Parth (parthkohli):

Looks like associativity :)

Parth (parthkohli):

\( \color{Black}{\Rightarrow a \times (b \times c) = (a \times b) \times c}\)